# Illusion or Algorithm? Investigating Memorization, Emergence, and Symbolic Processing in In-Context Learning

### v2

@@ 6.1 Method: Singular Residual Stream Direction Analysis
Here, we present a novel method to study the subdirection specialization process through Singular Residual Stream (Unembedding) Direction Analysis (SUDA). First, we perform a singular value decomposition (SVD) of the token unembedding ($W_{U}$) module [^9] We follow the notation and formulation of the residual stream proposed by Elhage et al. (2021).: $USV^{h}=\text{SVD}(W_{U})$. Then, we measure how well a single singular vector direction $V^{h}_{i}$ can perform the task. Specifically, we compute the logit of the answer token ($t_{\text{ans}}$) using the last layer’s residual output: $\ell(t_{\text{ans}})=V^{h}_{i}x_{-1}$.
Thus, for each pre-training Pythia checkpoint, we obtain a logit score for each singular component (i.e., each subdirection in the residual stream) for every input sample. For Pythia-12b [^10] We obtain the result in this section using the largest and most capable Pythia-12b model., the SVD process yields 5120 unique singular directions. As in our prior experiments, we feed the models 1000 samples across all tasks (lsc, lscg, wi, wc, tt, cf) using their default configurations [^11] We choose the settings where the task has reasonable complexity while the models can still obtain high performance. The particular settings are lsc: ||||P||||=5, ||||R||||=10; lsc: ||||P||||=5, ||||R||||=10, ||||G||||=2; wi: ||||S||||=5, $i$=1; and wc: ||||F||||=3, ||||L||||=2, ||||D||||=7. and take the average. We analyze the checkpoints timestep by timestep. We focus on two key statistics: (1) the maximum logit across all singular directions, and (2) how many “strong” directions exceed a given logit threshold. These two statistics, the max and spread, allow us to characterize how subdirections are utilized.
By observing these two statistics, we are especially interested in characterizing three key characteristics in this SUDA analysis: (1) whether individual subdirections can perform the tasks; (2) how many subdirections the model forms to perform a given task; and (3) how these characteristics develop throughout the pre-training process—are they related to the developmental characteristics we previously identified? We believe this novel method can be applied to a wide range of new tasks within the interpretability domain, and we welcome future work along with this direction.
6.2 The Formation of Subdirection Specialization Corresponds to ICL Generalization Ability
(a) Maximum logit value across subdirections at each training step. The curve shows a long-tailed shape, with the peak appearing at a very early stage ($\sim$1000 steps).
(b) Number of subdirections with logits crossing the threshold $\tau=0.2$. We can observe an early rise ($\sim$20,000 steps) followed by a gradual decline that remains above zero.
@@ 6.2 The Formation of Subdirection Specialization Corresponds to ICL Generalization Ability
Figure 11: The formation of subdirections in the residual stream follows the same trend (a,b) across all ICL tasks, even though their development differs drastically (c).
@@
These two trajectories of internal subdirection formation closely resemble how ICL generalization ability develops (as we explored in §5). First, recall that in-distribution generalization develops earlier, with the gap between in- and out-of-distribution generalization peaking around 1000 steps and then decreasing in a long-tailed manner. A similar pattern holds for the best-performing subdirection’s ability to reproduce the full model’s performance: the max logit peaks near 1000 steps, followed by a long-tailed decline. We believe this is not a coincidence. It is reasonable to assume that dominant subdirections capture the most frequent or structurally simple cases—those the model learns to handle correctly early in training. As shown earlier in the paper, ICL ability tends to appear first on common tokens before extending to rarer or more complex ones. In both cases, the metric reflects the performance of the single best-performing component, whether it is a subdirection or a performance difference across token frequency. The similarity in their trajectories is somewhat surprising but ultimately coherent.
@@
Second, recall that we observe a different curve when switching to Pearson correlation—a more comprehensive measure that captures nuanced levels of in- and out-of-distribution generalization, not just the extremes. We see an analogous trend in Figure 11(b) when extending the analysis to all subdirections, not just the maximum. The number of strong subdirections peaks at around 20,000—the same as in the Pearson correlation result—and is followed by a gradual decrease.
Thus, we hypothesize that the model’s internal ICL mechanisms are connected to its generalization ability. As discussed, there is evidence that these similarities and consistencies are not coincidental.
@@ 8 Conclusion
We also provide a more nuanced and accurate characterization on the development of ICL, tracing it back to the formation of internal mechanisms in LMs. In particular, we observe that different aspects of ICL competence develop at different rates, generally following the trend that “easier” aspects arise earlier and appear more rapid, while more complex aspects take longer to develop, presenting as gradual and predictable. These development are ultimately linked to the formation of internal mechanisms, which can be analyzed through allocation of subspaces in the residual stream. Therefore, our analysis goes beyond superficially probing model performance; instead, we propose a viable approach to bridging training dynamics analysis and mechanistic interpretability.
9 Acknowledgments
This work has been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing,” LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We would also like to thank the Early Career Research grant from the University of Bath.
@@ Appendix A Details of the Token Translation Task
Then, we manually examine the results generated by ChatGPT and narrow the list down to 200. Specifically, we prefer those that can be tokenised into a single token. The selected tokens are shown in Table 4 and 5.
@@ Appendix B Supplementary Experiment Results
Appendix C Supplementary Results: ICL Performance Strongly Correlates with Token Frequency

### v3

@@ 6.1 Method: Singular Residual Stream Direction Analysis
Here, we present a novel method to study the subspace specialization process through Singular Residual Stream (Unembedding) Direction Analysis (SUDA). First, we perform a singular value decomposition (SVD) of the token unembedding ($W_{U}$) module [^9] We follow the notation and formulation of the residual stream proposed by Elhage et al. (2021).: $USV^{h}=\text{SVD}(W_{U})$. Then, we measure how well a single singular vector direction $V^{h}_{i}$ can perform the task. Specifically, we compute the logit of the answer token ($t_{\text{ans}}$) using the last layer’s residual output: $\ell(t_{\text{ans}})=V^{h}_{i}x_{-1}$.
Thus, for each pre-training Pythia checkpoint, we obtain a logit score for each singular component for every input sample. For Pythia-12b [^10] We obtain the result in this section using the largest and most capable Pythia-12b model., the SVD process yields 5120 unique singular directions. As in our prior experiments, we feed the models 1000 samples across all tasks (lsc, lscg, wi, wc, tt, cf) using their default configurations [^11] We choose the settings where the task has reasonable complexity while the models can still obtain high performance. The particular settings are lsc: ||P||=5, ||R||=10; lsc: ||P||=5, ||R||=10, ||G||=2; wi: ||S||=5, ii=1; and wc: ||F||=3, ||L||=2, ||D||=7. and take the average. We analyze the checkpoints timestep by timestep. We focus on two key statistics: (1) the maximum logit across all singular directions, and (2) how many “strong” directions exceed a given logit threshold. These two statistics, the max and spread, allow us to characterize how subspaces are utilized.
By observing these two statistics, we are especially interested in characterizing three key characteristics in this SUDA analysis: (1) to what degree individual singular value directions contribute to a task; (2) how many singular value directions the model uses to perform a given task; and (3) how these characteristics develop throughout the pre-training process—are they related to the developmental characteristics we previously identified? We believe this novel method can be applied to a wide range of new tasks within the interpretability domain, and we welcome future work along with this direction.
6.2 The Formation of Subspace Specialization Corresponds to ICL Generalization Ability
(a) Maximum logit value across singular value directions at each training step. The curve shows a long-tailed shape, with the peak appearing at a very early stage ($\sim$1000 steps).
(b) Number of singular value directions with logits crossing the threshold $\tau=0.2$. We can observe an early rise ($\sim$20,000 steps) followed by a gradual decline that remains above zero.
@@ 6.2 The Formation of Subspace Specialization Corresponds to ICL Generalization Ability
Figure 11: The formation of subspaces in the residual stream follows the same trend (a,b) across all ICL tasks, even though their development differs drastically (c).
@@
These two trajectories of internal subspace formation closely resemble how ICL generalization ability develops (as we explored in §5). First, recall that in-distribution generalization develops earlier, with the gap between in- and out-of-distribution generalization peaking around 1000 steps and then decreasing in a long-tailed manner. A similar pattern holds for the best-performing subspace’s ability to reproduce the full model’s performance: the max logit peaks near 1000 steps, followed by a long-tailed decline. We believe this is not a coincidence. It is reasonable to assume that dominant singular value directions capture the most frequent or structurally simple cases—those the model learns to handle correctly early in training. As shown earlier in the paper, ICL ability tends to appear first on common tokens before extending to rarer or more complex ones. In both cases, the metric reflects the performance of the single best-performing component, whether it is a singular value direction or a performance difference across token frequency. The similarity in their trajectories is somewhat surprising but ultimately coherent.
@@
Second, recall that we observe a different curve when switching to Pearson correlation—a more comprehensive measure that captures nuanced levels of in- and out-of-distribution generalization, not just the extremes. We see an analogous trend in Figure 11(b) when extending the analysis to all singular value directions, not just the maximum. The number of strong singular value directions peaks at around 20,000—the same as in the Pearson correlation result—and is followed by a gradual decrease.
Table 4: Mean Overlap Matrix between Tasks.
Task lsc lscg wi wc $t$ cf CountryCapital
lsc - 0.632 0.597 0.593 0.303 0.224 0.207
lscg 0.632 - 0.608 0.578 0.305 0.215 0.206
wi 0.597 0.608 - 0.528 0.306 0.235 0.192
wc 0.593 0.578 0.528 - 0.297 0.204 0.231
$t$ 0.303 0.305 0.306 0.297 - 0.144 0.181
cf 0.224 0.215 0.235 0.204 0.144 - 0.096
Are these subspaces tied to ICL?
While these singular value directions are used by the model to perform ICL tasks, our analyses hitherto do not establish any exclusivity. To check whether certain singular value directions resonate with certain tasks, we look into the sharing of subspaces across tasks, defined by the average intersection over union (IoU) overlap throughout the training steps, defined as $\operatorname{IoU}(A,B)=\left(|V_{A}\cap V_{B}|\right)/\left(|V_{A}\cup V_{B}|\right)$, where $A,B$ are two tasks and $V_{A},V_{B}$ are their corresponding sets of singular value directions. As a control, we consider CountryCapital, a factual recall task that asks the model for the capital city of a given country, using the prompt template: The capital city of <country> is <city>. We use the underlying factual data from the cf task to generate this task. Unlike the other tasks, it does not require any non-trivial task learning from context.
In Table 4, we show the degree of singular value direction sharing among different tasks. We can see a high degree of sharing between the tasks devoid of any semantics, i.e., lsc, lscg, wc and wi. Requirement of pretrained knowledge introduces the usage of new subspace, and therefore the singular direction sharing between the non-semantic tasks and token translation (or counterfactual) decreases. For any of these ICL tasks, subspace sharing is lowest with the non-ICL factual recall task. Interestingly, the formation of singular value directions for the CountryCapital task follows a distinctly different trajectory compared to all the ICL tasks (see Figure 14 in Appendix E). We can conclude that: 1) development od ICL is exclusively tied to the formation of certain subspaces in the residual stream, 2) tasks that require pretrained knowledge can be associated with distinctly separate subspaces, and 3) formation (over pre-training) of these two classes of subspaces are distinctly different. We hypothesize that, these subspaces arise from the distinct subspaces of the low-rank attention heads: certain heads specialize in moving information from the context (Olsson et al., 2022) whereas other heads act as movers of pretrained knowledge (Lv et al., 2024; Singh et al., 2024). We leave the training dynamics of the attention head-specific subspaces and their alignment as an important future work.
@@ 8 Conclusion
We also provide a more nuanced and accurate characterization on the development of ICL, tracing it back to the formation of internal mechanisms in LMs. In particular, we observe that different aspects of ICL competence develop at different rates, generally following the trend that “easier” aspects arise earlier and appear more rapid, while more complex aspects take longer to develop, presenting as gradual and predictable. These development are ultimately linked to the formation of internal mechanisms, which can be analyzed through the allocation of subspaces in the residual stream. Therefore, our analysis goes beyond superficially probing model performance; instead, we propose a viable approach to bridging training dynamics analysis and mechanistic interpretability.
@@ Appendix A Details of the Token Translation Task
Then, we manually examine the results generated by ChatGPT and narrow the list down to 200. Specifically, we prefer those that can be tokenised into a single token. The selected tokens are shown in Table 5 and 6.
@@ Appendix B Supplementary Experiment Results
Appendix C Supplementary Results I: ICL Performance Strongly Correlates with Token Frequency – Llama Models
@@
Appendix D Supplementary Results II: ICL Performance Strongly Correlates with Token Frequency Across Tasks
In this section, we list the token frequency experiments for other literal-based tasks. As shown in Figure 13 and Table 12, we see that the same finding persists: ICL performance strongly correlates with token frequency across tasks.
(a) lscg
(b) wc
(c) wi
Figure 13: LM performance on other literal-based tasks (lscg, wc, and wi) also deteriorates as tokens are sampled from ranges with larger indices. Line plot colours correspond to model size, following the same colour scheme as Figure 1.
Model Size 410m 1b 1.4b 2.8b 6.9b 12b
Task: lscg
Pearson $r$ -0.742 -0.731 -0.876 -0.842 -0.588 -0.840
pp-value 1.040E-09 2.531E-09 1.829E-16 3.401E-14 8.947E-06 4.357E-14
Task: wc
Pearson $r$ -0.501 -0.701 -0.763 -0.736 -0.775 -0.852
pp-value 2.482E-04 2.010E-08 1.931E-10 1.623E-09 6.412E-11 8.612E-15
Task: wi
Pearson $r$ -0.716 -0.639 -0.684 -0.814 -0.820 -0.805
pp-value 7.119E-09 7.829E-07 6.136E-08 1.199E-12 5.551E-13 3.043E-12
Appendix E SUDA Analysis for Factual Tasks
Figure 14: Threshold dynamics of singular value directions in the CityCapital task, showing a distinct trajectory compared to ICL tasks.
To examine whether the observed subspace formation trend generalises beyond ICL, we applied the same SUDA analysis to a factual recall task, CityCapital, which requires retrieving the capital city given a country name. Interestingly, the formation of singular value directions in this task follows a distinctly different trajectory from that observed in the ICL tasks, with no clear rise of the dominant components that characterise in-context learning. This suggests that the progressive alignment of singular directions may be a property specific to ICL rather than a universal feature of model adaptation. A more systematic investigation across diverse factual and reasoning tasks is left for future work.